OpenStudy (anonymous):

Solve the following system of linear equations: -eq is posted in reply- If the system has infinitely many solutions, your answer may use expressions involving the parameters r, s, and t.

OpenStudy (anonymous):

4

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

don't know

OpenStudy (anonymous):

\[ x _{1}+5x _{2}−5x _{3} = 5 .\]

OpenStudy (anonymous):

I'll post in just one second.

OpenStudy (anonymous):

First of all, let's be clear that this system has infinitely many solutions. Now I have a question: are you supposed to be finding the vectors that span the solution space (or whatever the correct linear algebra terminology is)? Or is it basically like simplifying the system as much as possible (reduced row echelon form) and then relating the variables? Because if it's the latter I think it'd be as simple as saying x1=5-5s+5t where s=x2 and t=x3.

OpenStudy (anonymous):

Yes, you are suppose to simplify this to R.E.F and once you have it all expanded. it would have each x1, x2, x3= to a variable. I am having difficulties just putting it into R.E.F

OpenStudy (anonymous):

Well good news: it's already in REF. lol

OpenStudy (anonymous):

I'll post the matrix in just a sec.

OpenStudy (anonymous):

lol, you are amazing

OpenStudy (anonymous):

Which phrase best defines the term direct object? A. an adverb that modifies an action verb B. a word connected to the subject by a linking verb C. a noun, pronoun, or word group that receives the action of the verb D. a noun or pronoun before an action verb

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

this is a math forum. please go learn english in another field

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 5 & -5 \\ 0 & 0 & 0\\ 0 & 0 & 0\ &\end{matrix}\right]*x=\left[\begin{matrix}5 \\ 0 \\ 0\end{matrix}\right]\] where x is a vector.

OpenStudy (anonymous):

Notice that all rows start with leading ones, which means this is in REF.

OpenStudy (anonymous):

Oh ya! that is right.

OpenStudy (anonymous):

Do you care about the other thing I was talking about at all or that too complicated?

OpenStudy (anonymous):

I would really like to know how would you solve for x1. x2. and x3. I understand now that i can physically see in a R.E.F, but would like how to continue from there

OpenStudy (anonymous):

Well solve for x1: x1=-5x2+5x3+5

OpenStudy (anonymous):

Now let x2 and x3 be equal to different parametric values: let x2=s and x3=t for instance. So x1=-5x2+5x3+5=-5s+5t+5

OpenStudy (anonymous):

ohhhh. so would the x2= be just corresponding to s and x3= t

OpenStudy (anonymous):

It is worth noting that you could solve for any of the 3 xn but typically we solve for the lowest aka the one corresponding to the leading 1 in REF. And yeah pretty much.

OpenStudy (anonymous):

wow, this is much easier than what I assumed before

OpenStudy (anonymous):

You are amazing! :) thank you so much.

OpenStudy (anonymous):

I'll go ahead and post the vectors that span the solution space just for the heck of it. Just one second.

OpenStudy (anonymous):

I have another question too regarding about this type of question. I hope you can help me out too! :)

OpenStudy (anonymous):

\[x=\left[\begin{matrix} x _{1} \\ x _{2}\\ x _{3}\\ \end{matrix}\right]=\left[\begin{matrix} -5 \\ 1\\ 0\\ \end{matrix}\right]s+\left[\begin{matrix}5\\0\\1\\ \end{matrix}\right]t+\left[\begin{matrix} 5 \\0\\0\\ \end{matrix}\right]\]

OpenStudy (anonymous):

I'll do my best. Go for it.

OpenStudy (anonymous):

in another forum? or here?

OpenStudy (anonymous):

Typically when you're looking for the span of a space though you have Ax=0 instead of Ax=b as we have here and that (5,0,0) wouldn't be there.

OpenStudy (anonymous):

You can put it here if you want.

OpenStudy (anonymous):

Sweet. Thank you so much!

OpenStudy (anonymous):

Determine the values of a for which the following system of linear equations has no solutions, a unique solution, or infinitely many solutions. 5x1+5x2+5x3 = 10 −3x1−2x2+ax3 = −4 x1+4x2+4x3 = 9

OpenStudy (anonymous):

In that case, I am trying to find how this would be an unique solution, infintely many, and no solutions

OpenStudy (anonymous):

\[\left[\begin{matrix}5 & 5 & 5 \\ -3 & -2 & a\\ 1 & 4 & 4 \end{matrix}\right]x=\left[\begin{matrix} 10 \\ -4\\ 9\end{matrix}\right]\]

OpenStudy (anonymous):

One way is to put it in RREF and then solve for a to make those conditions true.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Excuse me REF not RREF.

OpenStudy (anonymous):

lol :)

OpenStudy (anonymous):

Ok I'm getting\[\left[\begin{matrix}1 & 4 & 4 \\ 0 & 1 & 1\\ 0 & 0 & a+2\end{matrix}\right]x+\left[\begin{matrix} 9 \\ 7/3\\ 23-70/3\end{matrix}\right]\]

OpenStudy (anonymous):

ok, i got that too despite fractions

OpenStudy (anonymous):

So if it were going to be infinitely many solutions, you would find that you have less than 3 independent equations. So you have only 1 or 2 leading ones and that the last row is all 0. So we can let a+2=0 which means a=-2 for infinitely many solutions.

OpenStudy (anonymous):

yep, got that too

OpenStudy (anonymous):

Oh I know. A better way is to then take the determinant of the matrix.

OpenStudy (anonymous):

a determiniant?

OpenStudy (anonymous):

Have you not done that before?

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

Ok we won't do that then lol. Ok back to what we were doing then.

OpenStudy (anonymous):

Ok recapping: we have that if a=-2, 0=23-70/3 which cannot be. This is the no solution case (not the infinitely many solutions case like I mistakenly said earlier). For infinitely many solutions we need 2 more more of the equations to be the same.

OpenStudy (anonymous):

interesting

OpenStudy (anonymous):

Yeah sorry about that XD.

OpenStudy (anonymous):

haha no problem.

OpenStudy (anonymous):

Grrr I see I messed up the REF let me change that too. My arithmetic is terrible today.

OpenStudy (anonymous):

hahah same with mine!

OpenStudy (anonymous):

this course really defines how good i am at arthmetic

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 1 & 1\\ 0 & 1 & 1\\ 0& 0 & a+2\end{matrix}\right]x=\left[\begin{matrix} 2 \\ 7/4\\ 1/4\end{matrix}\right]\] Is this the REF you got?

OpenStudy (anonymous):

This however, won't change the fact that a=-2 makes this have no solution since it's still 0=something which can't happen.

OpenStudy (anonymous):

yep, i corrected mine too.my row 3 was messed

OpenStudy (anonymous):

Anyway, If we subtract the 2nd equation from the first:\[\left[\begin{matrix}1 & 0 & 0 \\ 0 & 1 & 1\\ 0 & 0 & a+2\end{matrix}\right]x=\left[\begin{matrix} 1/4 \\7/4\\ 1/4\end{matrix}\right]\]

OpenStudy (anonymous):

Nah nevermind that's not so useful, after all.

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

I'm not seeing the way to make this have infinitely many solutions. There's something I'm missing or there is a unique for solution for any value of a that isn't a=-2. I'll keep looking and get back to you in a bit.

OpenStudy (anonymous):

arlighty

OpenStudy (anonymous):

Here is what I got for RREF btw:\[\left[\begin{matrix}1 & 0 & 0 \\ 0 & 1 & 0\\ 0 & 0 & 1\end{matrix}\right]x=\left[\begin{matrix} 1/4 \\ 7/4-\frac{ 1 }{ 4(a+2) }\\ \frac{ 1 }{ 4(a+2) }\ \end{matrix}\right]\]

OpenStudy (anonymous):

woah. thats crazy

OpenStudy (anonymous):

Actually I'm fairly certain the matrix can't be both no solutions and infinitely many solutions. Think about it in 2D: What does the graph look like?